# Models of Set Theory: Extensions and Dead-ends

### v5

@@ Abstract.
This paper is a contribution to the study of extensions of arbitrary models of $\mathrm{ZF}$ (Zermelo-Fraenkel set theory), with no regard to countability or well-foundedness of the models involved. Our main results include the theorems below; in Theorems A and B, $\mathcal{N}$ is said to be a conservative elementary extension of $\mathcal{M}$ if $\mathcal{N}$ elementarily extends $\mathcal{M}$, and the intersection of every $\mathcal{N}$-definable set with $M$ is $\mathcal{M}$-definable (parameters allowed). In Theorem B, ZFC is the result of augmenting ZF with the axiom of choice.
Theorem A. Every model $\mathcal{M}$ of $\mathrm{ZF+}\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right)$ has a conservative elementary extension $\mathcal{N}$ that contains an ordinal above all of the ordinals of $\mathcal{M}$.
Theorem B. If $\mathcal{N}$ is a conservative elementary extension of a model $\mathcal{M}$ of $\mathrm{ZFC}$, and $\mathcal{N}$ has the same natural numbers as $\mathcal{M}$, then $\mathcal{M}$ is cofinal in $\mathcal{N}$.
Theorem C. Every consistent extension of $\mathrm{ZFC}$ has a model $\mathcal{M}$ of power $\aleph_{1}$ such that $\mathcal{M}$ has no proper end extension to a model of $\mathrm{ZF}$.
@@ 1. INTRODUCTION
This paper concerns the model theory of Zermelo-Fraenkel ($\mathrm{ZF}$) set theory. More specifically, it concerns the types of extensions that are possible for arbitrary models of $\mathrm{ZF}$ (as opposed to countable or well-founded ones), a topic that I have worked on intermittently for over four decades. Indeed, Theorem C of the abstract answers the following question that was initially posed in my 1984 doctoral dissertation [En-1, Question 1.1.12]:
@@
Note that assuming that strongly inaccessible cardinals are cofinal in the class of ordinals, every well-founded model of $\mathrm{ZF}$ can be properly end extended to a model of $\mathrm{ZF}$. The motivation for the above question emerges from the comparative study of the model theory of $\mathrm{ZF}$ and $\mathrm{PA}$ (Peano Arithmetic). Prima facie $\mathrm{ZF}$ and $\mathrm{PA}$ are unrelated; after all $\mathrm{PA}$ axiomatizes basic intuitions concerning the familiar arithmetical structure of natural numbers, whereas $\mathrm{ZF}$ codifies the laws governing Cantor’s mysterious universe of sets, a vastly more complex structure. However, thanks to a clever coding idea, introduced first by Ackermann [A], $\mathrm{ZF}$ and $\mathrm{PA}$ turn out to be intimately connected at a formal level: $\mathrm{PA}$ is bi-interpretable [^2] The role of TC was elucidated in [ESV] by showing that TC cannot be dropped in this bi-interpretability result. The details of this bi-interpretability result are worked out by Kaye and Wong in [KW]. Their work shows that indeed the two theories are definitionally equivalent (aka synonymous), i.e., they have a common definitional extension. with the theory $\mathrm{ZF}_{\mathrm{fin}}+\mathrm{TC}$, where $\mathrm{ZF}_{\mathrm{fin}}$ is obtained from $\mathrm{ZF}$ by replacing the axiom of infinity by its negation, and $\mathrm{TC}$ expresses “the transitive closure of every set exists”. In this light it is natural to compare and contrast the model theoretic behavior of models of $\mathrm{ZF}$ and $\mathrm{PA}$ in order to elucidate the role of the axiom of infinity in Zermelo-Fraenkel set theory. A central result in the model theory of PA is the MacDowell-Specker Theorem [MS], and its refinement (independently) by Phillips [Ph] and Gaifman [G-2], that states that every model of PA (of any cardinality) has a proper conservative elementary end extension. Using the aforementioned bi-interpretation between $\mathrm{PA}$ and $\mathrm{ZF}$, the refinement of the MacDowell-Specker Theorem is equivalent to:
@@
In the above, $\mathcal{N}$ is said to be a conservative elementary extension of $\mathcal{M}$ if $\mathcal{N}$ elementarily extends $\mathcal{M}$, and the intersection of every $\mathcal{N}$-definable set with $M$ is $\mathcal{M}$-definable (parameters allowed). It is known that Theorem $\triangledown$ becomes false if $\mathrm{ZF}_{\mathrm{fin}}+\mathrm{TC}$ is replaced by $\mathrm{ZF}$, but the answer to Question $\heartsuit$ (Theorem 5.18) remained open until now. [^3] By a remarkable theorem of Keisler and Morley [KS] every model of $\mathrm{ZF}$ of countable cofinality has arbitrarily large elementary end extensions, however, as shown by Keisler and Silver [KS] not every model of $\mathrm{ZF}$ has an elementary end extension. For an exposition of the general framework for the analogues of the MacDowell-Specker Theorem for set theory, see [En-5].
@@
Section 2 is devoted to preliminaries needed for Sections 3, 4, and 5 in which the main results are presented. Section 3 contains the proof of Theorem A of the abstract (and closely related material). Section 4 contains results on faithful extensions (a generalization of conservative extensions), including Theorem 4.1 that yields Theorem B of the abstract (as Corollary 4.2). A key result in Section 4 is Theorem 4.4 which is used in Section 5 in conjunction with other machinery to establish Theorem C of the abstract. Section 6 is an Addendum and Corrigendum to my aforementioned 1987 paper [En-3]. The Appendix presents the proof of a theorem due to Rubin, Shelah, and Schmerl that plays a key role in the proof of Theorem C of the abstract.
The main results of the paper were presented online in February 2025 at the CUNY set theory seminar. I am grateful to Vika Gitman and Gunter Fuchs and other participants of the seminar for helpful feedback. I am also grateful to Corey Switzer for illuminating discussions about the forcing argument used in the proof of Step 2 of Stage 2 of Theorem A.1 (in the Appendix).
@@ 2. PRELIMINARIES
(a) $\mathcal{L}_{\mathrm{set}}=\{=,\in\}$ is the usual language of $\mathrm{ZF}$-set theory. We treat $\mathrm{ZF}$ as being axiomatized as usual, except that instead of including the scheme of replacement among the axioms of $ZF$, we include the schemes of separation and collection, as in [CK, Appendix A]. Thus, in our setup the axioms of Zermelo set theory $\mathrm{Z}$ are obtained by removing the scheme of collection from the axioms of $\mathrm{ZF}$. More generally, for $\mathcal{L}\supseteq\mathcal{L}_{\mathrm{set}}$ we construe $\mathrm{ZF}(\mathcal{L})$ to be the natural extension of $\mathrm{ZF}$ in which the schemes of separation and collection are extended to $\mathcal{L}$-formulae; similarly we will use $\mathrm{Z}(\mathcal{L})$ to denote Zermelo set theory over $\mathcal{L}$, i.e., as the result of extending $\mathrm{Z}$ with the $\mathcal{L}$-separation scheme $\mathrm{Sep}(\mathcal{L})$. Officially speaking, $\mathrm{Sep}(\mathcal{L})$ consists of the universal closures of $\mathcal{L}$-formulae of the form:
$\forall v\exists w\forall x(x\in w\longleftrightarrow x\in v\wedge\varphi(x,\overline{y})).$
@@
Thus $\mathrm{ZF}(\mathcal{L})$ is the result of augmenting $\mathrm{Z}(\mathcal{L})$ with the $\mathcal{L}$-collection scheme $\mathrm{Coll}(\mathcal{L})$, which consists of the universal closures of $\mathcal{L}$-formulae of the form:
$\left(\forall x\in v\text{ }\exists y\text{\ }\varphi(x,y,\overline{z})\right)\rightarrow\left(\exists w\text{ }\forall x\in v\text{ }\exists y\in w\text{ }\varphi(x,y,\overline{z})\right).$
When X is a new predicate, we write $\mathcal{L}_{\mathrm{set}}(\mathrm{X})$ instead of $\mathcal{L}_{\mathrm{set}}\cup\{\mathrm{X\}}$, and similarly we write $\mathrm{ZF}(\mathrm{X})$, $\mathrm{Sep}(X)$, $\mathrm{Coll}(X)$, etc. instead of $\mathrm{ZF}(\mathcal{L})$, $\mathrm{Sep}(\mathcal{L})$, $\mathrm{Coll}(\mathcal{L})$, etc. (respectively) when $\mathcal{L}=\mathcal{L}_{\mathrm{set}}(\mathrm{X})$.
(b) For $n\in\omega$, we employ the common notation ($\Sigma_{n},$ $\Pi_{n},$ $\Delta_{n}$) for the Levy hierarchy of $\mathcal{L}_{\mathrm{set}}$-formulae, as in the standard references in modern set theory by Kunen [Ku], Jech [J] and Kanamori [Kan]. In particular, $\Delta_{0}=\Sigma_{0}=\Pi_{0}$ corresponds to the collections of $\mathcal{L}_{\mathrm{set}}$-formulae all of whose quantifiers are bounded. In addition, we will also consider the Takahashi hierarchy of formulae ($\Sigma_{n}^{\mathcal{P}},$ $\Pi_{n}^{\mathcal{P}},$ $\Delta_{n}^{\mathcal{P}}$) (introduced in [Takahashi], and further studied by Mathias’ [M]), where $\Delta_{0}^{\mathcal{P}}$ is the smallest class of $\mathcal{L}_{\mathrm{set}}$-formulae that contains all atomic formulae, and is closed both under Boolean connectives and under quantification in the form $Qx\in y$ and $Qx\subseteq y$ where $x$ and $y$ are distinct variables, and $Q$ is $\exists$ or $\forall$. The classes $\Sigma_{1}^{\mathcal{P}},\Pi_{1}^{\mathcal{P}}$, etc. are defined inductively from the class $\Delta_{0}^{\mathcal{P}}$ in the same way that the formula-classes $\Sigma_{1},\Pi_{1}$, etc. are defined from $\Delta_{0}$-formulae.
(c) For $n\in\omega$. and $\mathcal{L}\supseteq\mathcal{L}_{\mathrm{set}},$ the Levy hierarchy can be naturally extended to $\mathcal{L}$-formulae ($\Sigma_{n}(\mathcal{L}),$ $\Pi_{n}(\mathcal{L}),$ $\Delta_{n}(\mathcal{L})$), where $\Delta_{0}(\mathcal{L})$ is the smallest family of $\mathcal{L}$-formulae that contains all atomic $\mathcal{L}$-formulae and is closed under Boolean operations and bounded quantification.
(d) $\mathrm{KPR}$ (Kripke-Platek with ranks) is the subtheory of $\mathrm{ZF}$ whose axioms consist of KP plus an axiom that asserts that for all ordinals $\alpha,$$\{x:\mathrm{rank}(x)<\alpha\}$ exists (where $\mathrm{rank}(x)$ is the usual rank function on sets). Following recent practice (initiated by Mathias [M]), the foundation scheme of $\mathrm{KP}$ is only limited to $\Pi_{1}$-formulae; thus in this formulation $\mathrm{KP}$ can prove the scheme of $\in$-induction for $\Sigma_{1}$-formulae. In contrast to Barwise’s KP in [Bar], which includes the full scheme of foundation, our version of KP is finitely axiomatizable. Thus $\mathrm{KPR}$ is finitely axiomatizable.
(e) For $\mathcal{L}\supseteq\mathcal{L}_{\mathrm{set}}$, the dependent choice scheme, denoted $\Pi_{\infty}^{1}$-$\mathrm{DC(}\mathcal{L}\mathrm{)}$, consists of $\mathcal{L}$-sentences of the following form:
$\forall z\ \left[\forall x\exists y\ \varphi(x,y,z)\longrightarrow\forall x\exists f\ \left(f:\omega\rightarrow\mathrm{V},\ f(0)=x,\mathrm{and}\ \forall n\in\omega\ \varphi(f(n),f(n+1),z)\right)\right].$
In the presence of $\mathrm{ZF(\mathcal{L}),}$ thanks to the Reflection Theorem, $\Pi_{\infty}^{1}$-$\mathrm{DC(\mathcal{L})}$ is equivalent to the single sentence $\mathrm{DC}$ below:
@@
Note that $\mathrm{DC}$ is provable in $\mathrm{ZFC}$.
@@
(a) We follow the convention of using $M$, $M^{\ast},$ $M_{0}$, etc. to denote (respectively) the universes of structures $\mathcal{M}$, $\mathcal{M}^{\ast},$ $\mathcal{M}_{0},$ etc. We denote the membership relation of $\mathcal{M}$ by $\in^{\mathcal{M}}$; thus an $\mathcal{L}_{\mathrm{set}}$-structure $\mathcal{M}$ is of the form $(M,\in^{\mathcal{M}})$.
(b) $\mathrm{Ord}^{\mathcal{M}}$ is the class of “ordinals” of $\mathcal{M}$, i.e.,
@@
where $\mathrm{Ord}(x)$ expresses “$x$ is transitive and is well-ordered by $\in$”. More generally, for a formula $\varphi(\overline{x})$ with parameters from $M$, where $\overline{x}=\left(x_{1},\cdot\cdot\cdot,x_{k}\right)$, we write $\varphi^{\mathcal{M}}$ for $\left\{\overline{m}\in M^{k}:\mathcal{M\models\varphi}\left(m_{1},\cdot\cdot\cdot,m_{k}\right)\right\}.$ A subset $D$ of $M^{k}$ is $\mathcal{M}$-definable if it is of the form $\varphi^{\mathcal{M}}$ for some choice of $\varphi.$ We write $\mathbb{\omega}^{\mathcal{M}}$ for the set of finite ordinals (i.e., natural numbers) of $\mathcal{M}$, and $\mathbb{\omega}$ for the set of finite ordinals in the real world, whose members we refer to as metatheoretic natural numbers. $\mathcal{M}$ is said to be $\omega$-standard if $\left(\mathbb{\omega},\mathbb{\in}\right)^{\mathcal{M}}\cong\left(\mathbb{\omega},\mathbb{\in}\right).$ For $\alpha\in\mathrm{Ord}^{\mathcal{M}}$ we often use $\mathcal{M}_{\alpha}$ to denote the substructure of $\mathcal{M}$ whose universe is $\mathrm{V}_{\alpha}^{\mathcal{M}},$ here $\mathrm{V}_{\alpha}$ is defined as usual within $\mathrm{ZF}$ as $\{x:\mathrm{rank}(x)<\alpha\}$, where $\mathrm{rank}(x)=\sup\{\mathrm{rank}(y)+1:y\in x\}.$

### v6

@@ Abstract.
This paper is a contribution to the study of extensions of arbitrary models of $\mathsf{ZF}$ (Zermelo-Fraenkel set theory), with no regard to countability or well-foundedness of the models involved. Our main results include the theorems below; in Theorems A and B, $\mathcal{N}$ is said to be a conservative elementary extension of $\mathcal{M}$ if $\mathcal{N}$ elementarily extends $\mathcal{M}$, and the intersection of every $\mathcal{N}$-definable set with the universe of $\mathcal{M}$ is $\mathcal{M}$-definable (parameters allowed). In Theorem B, $\mathsf{ZFC}$ is the result of augmenting $\mathsf{ZF}$ with the axiom of choice.
Theorem A. Every model $\mathcal{M}$ of $\mathsf{ZF}+\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right)$ has a conservative elementary extension $\mathcal{N}$ that contains an ordinal above all of the ordinals of $\mathcal{M}$.
Theorem B. If $\mathcal{N}$ is a conservative elementary extension of a model $\mathcal{M}$ of $\mathsf{ZFC}$, and $\mathcal{N}$ has the same natural numbers as $\mathcal{M}$, then $\mathcal{M}$ is cofinal in $\mathcal{N}$.
Theorem C. Every consistent extension of $\mathrm{ZFC}$ has a model $\mathcal{M}$ of power $\aleph_{1}$ such that $\mathcal{M}$ has no proper end extension to a model of $\mathsf{ZF}$.
@@ 1. INTRODUCTION
This paper concerns the model theory of Zermelo-Fraenkel ($\mathsf{ZF}$) set theory. More specifically, it is about various types of extensions of arbitrary models of $\mathsf{ZF}$, as opposed to countable or well-founded ones. This is a topic that I have worked on intermittently for over four decades. Indeed, Theorem C of the abstract answers the following question that was initially posed in my 1984 doctoral dissertation [En-1, Question 1.1.12]:
@@
Note that if there is a proper class of strongly inaccessible cardinals, then every well-founded model of $\mathsf{ZF}$ can be properly end extended to a model of $\mathsf{ZF}$. Furthermore, by a well-known result due to Keisler and Morley [KM], every model of $\mathsf{ZF}$ whose class of ordinals has countable cofinality has a proper elementary end extension. The motivation for the above question emerges from the comparative study of the model theory of $\mathsf{ZF}$ and $\mathsf{PA}$ (Peano Arithmetic). Prima facie, $\mathsf{ZF}$ and $\mathsf{PA}$ are unrelated; after all $\mathsf{PA}$ axiomatizes basic intuitions concerning the familiar arithmetical structure of natural numbers, whereas $\mathsf{ZF}$ codifies the laws governing Cantor’s mysterious universe of sets, a vastly more complex structure. However, thanks to a clever coding idea, introduced first by Ackermann [A], $\mathsf{ZF}$ and $\mathsf{PA}$ turn out to be intimately connected at a formal level: $\mathsf{PA}$ is bi-interpretable with the theory $\mathsf{ZF}_{\mathrm{fin}}+\mathsf{TC}$, where $\mathsf{ZF}_{\mathrm{fin}}$ is obtained from $\mathsf{ZF}$ by replacing the axiom of infinity by its negation, and $\mathsf{TC}$ expresses “the transitive closure of every set exists”. [^2] The role of $\mathsf{TC}$ was elucidated in [ESV] by showing that $\mathsf{TC}$ cannot be dropped in this bi-interpretability result. The details of this bi-interpretability result are worked out by Kaye and Wong in [KW]. Their work shows that indeed the two theories are definitionally equivalent (aka synonymous), i.e., they have a common definitional extension. In this light it is natural to compare and contrast the model theoretic behavior of models of $\mathsf{ZF}$ and $\mathsf{PA}$ in order to elucidate the role of the axiom of infinity in Zermelo-Fraenkel set theory. A central result in the model theory of $\mathsf{PA}$ is the MacDowell-Specker Theorem [MS], and its refinement (independently) by Phillips [P] and Gaifman [G-2], that states that every model of PA (of any cardinality) has a proper conservative elementary end extension. Using the aforementioned bi-interpretation between $\mathsf{PA}$ and $\mathsf{ZF}$, the refinement of the MacDowell-Specker Theorem is equivalent to:
@@
In the above, $\mathcal{N}$ is said to be a conservative elementary extension of $\mathcal{M}$ if $\mathcal{N}$ elementarily extends $\mathcal{M}$, and the intersection of every $\mathcal{N}$-definable set with $M$ is $\mathcal{M}$-definable (parameters allowed). It is known that Theorem $\triangledown$ becomes false if $\mathsf{ZF}_{\mathrm{fin}}+\mathsf{TC}$ is replaced by $\mathsf{ZF}$, but the answer to Question $\heartsuit$ (Theorem 5.18) remained open until now. [^3] As shown by Keisler and Silver [KS], if $\kappa$ is the first inaccessible cardinal, then $\left(\mathrm{V}_{\kappa},\in\right)$ does not have a proper elementary end extension. More generally, the results of Kaufmann [Kan] and the author [En-2] showed every consistent extension of $\mathsf{ZFC}$ has a model of power $\aleph_{1}$ that has no proper rank extension to another model of $\mathsf{ZFC}$. For an exposition of the general framework for the analogues of the MacDowell-Specker Theorem for set theory, see [En-5].
@@
Section 2 is devoted to preliminaries needed for Sections 3, 4, and 5 in which the main results are presented. Section 3 contains the proof of Theorem A of the abstract (and closely related material). Section 4 contains results on faithful extensions (a generalization of conservative extensions), including Theorem 4.1 that yields Theorem B of the abstract (as Corollary 4.2). A key result in Section 4 is Theorem 4.4 which is used in Section 5 in conjunction with other machinery to establish Theorem C of the abstract. Section 6 is an Addendum and Corrigendum to my aforementioned 1987 paper [En-3]. The Appendix presents the proof of a theorem due to the collective effort of Rubin, Shelah, and Schmerl that plays a key role in the proof of Theorem C of the abstract.
The main results of the paper were presented online in February 2025 at the CUNY set theory seminar. I am grateful to Vika Gitman and Gunter Fuchs and other participants of the seminar for helpful feedback. Thanks also to Corey Switzer for illuminating discussions about the forcing argument used in the proof of Theorem A.1 (in the Appendix), and to Junhong Chen for catching notational inconsistencies of a previous draft. I am especially indebted to the anonymous referee for meticulous and extensive feedback that was instrumental in shaping the paper in its current form.
@@ 2. PRELIMINARIES
    $\mathcal{L}_{\mathrm{set}}=\{=,\in\}$ is the language of $\mathsf{ZF}$-set theory. We treat $\mathsf{ZF}$ as being axiomatized as usual, except that instead of including the scheme of replacement among the axioms of $\mathsf{ZF}$, we include the schemes of separation and collection, as in [CK, Appendix A]. Thus, in our setup the axioms of Zermelo set theory $\mathsf{Z}$ are obtained by removing the scheme of collection from the axioms of $\mathsf{ZF}$. More generally, for $\mathcal{L}\supseteq\mathcal{L}_{\mathrm{set}}$ we construe $\mathsf{ZF}(\mathcal{L})$ to be the natural extension of $\mathsf{ZF}$ in which the schemes of separation and collection are extended to $\mathcal{L}$-formulae; similarly we will use $\mathsf{Z}(\mathcal{L})$ to denote Zermelo set theory over $\mathcal{L}$, i.e., as the result of extending $\mathsf{Z}$ with the $\mathcal{L}$-separation scheme $\mathsf{Sep}(\mathcal{L})$. Officially speaking, $\mathsf{Sep}(\mathcal{L})$ consists of the universal closures of $\mathcal{L}$-formulae of the form
    $\forall v\exists w\forall x(x\in w\longleftrightarrow x\in v\wedge\varphi(x,\vec{y})).$
@@
    Thus $\mathsf{ZF}(\mathcal{L})$ is the result of augmenting $\mathsf{Z}(\mathcal{L})$ with the $\mathcal{L}$-collection scheme $\mathsf{Coll}(\mathcal{L})$, which consists of the universal closures of $\mathcal{L}$-formulae of the form:
@@
    When X is a new predicate, we write $\mathcal{L}_{\mathrm{set}}(\mathrm{X})$ instead of $\mathcal{L}_{\mathrm{set}}\cup\{\mathrm{X\}}$, and similarly we write $\mathsf{ZF}(\mathrm{X})$, $\mathsf{Sep}(\mathrm{X})$, $\mathsf{Coll}(\mathrm{X})$, etc. instead of $\mathsf{ZF}(\mathcal{L})$, $\mathsf{Sep}(\mathcal{L})$, $\mathsf{Coll}(\mathcal{L})$, etc. (respectively) when $\mathcal{L}=\mathcal{L}_{\mathrm{set}}(\mathrm{X})$.
  2. (b)
    For $n\in\omega$, we employ the common notation ($\Sigma_{n},$ $\Pi_{n},$ $\Delta_{n}$) for the Levy hierarchy of $\mathcal{L}_{\mathrm{set}}$-formulae, as in the standard references in modern set theory by Kunen [Ku], Jech [J] and Kanamori [Kan]. In particular, $\Delta_{0}=\Sigma_{0}=\Pi_{0}$ corresponds to the collections of $\mathcal{L}_{\mathrm{set}}$-formulae all of whose quantifiers are bounded. In addition, we will also consider the Takahashi hierarchy of formulae ($\Sigma_{n}^{\mathcal{P}},$ $\Pi_{n}^{\mathcal{P}},$ $\Delta_{n}^{\mathcal{P}}$) (introduced in [T], and further studied by Mathias’ [M-1]), where $\Delta_{0}^{\mathcal{P}}$ is the smallest class of $\mathcal{L}_{\mathrm{set}}$-formulae that contains all atomic formulae, and is closed both under Boolean connectives and under quantification in the form $Qx\in y$ and $Qx\subseteq y$ where $x$ and $y$ are distinct variables, and $Q$ is $\exists$ or $\forall$. The classes $\Sigma_{1}^{\mathcal{P}},\Pi_{1}^{\mathcal{P}}$, etc. are defined inductively from the class $\Delta_{0}^{\mathcal{P}}$ in the same way that the formula-classes $\Sigma_{1},\Pi_{1}$, etc. are defined from $\Delta_{0}$-formulae.
  3. (c)
    For $n\in\omega$. and $\mathcal{L}\supseteq\mathcal{L}_{\mathrm{set}},$ the Levy hierarchy can be naturally extended to $\mathcal{L}$-formulae ($\Sigma_{n}(\mathcal{L}),$ $\Pi_{n}(\mathcal{L}),$ $\Delta_{n}(\mathcal{L})$), where $\Delta_{0}(\mathcal{L})$ is the smallest family of $\mathcal{L}$-formulae that contains all atomic $\mathcal{L}$-formulae and is closed under Boolean operations and bounded quantification.
  4. (d)
    $\mathsf{KPR}$ (Kripke-Platek with ranks) is the subtheory of $\mathsf{ZF}$ whose axioms consist of $\mathsf{KP}$ plus an axiom that asserts that for all ordinals $\alpha,$ $\{x:\mathrm{rank}(x)<\alpha\}$ exists (where $\mathrm{rank}(x)$ is the usual rank function on sets). Following recent practice (initiated by Mathias [M-1]), the foundation scheme of $\mathsf{KP}$ is only limited to $\Pi_{1}$-formulae; thus in this formulation $\mathsf{KP}$ can prove the scheme of $\in$-induction for $\Sigma_{1}$-formulae. In contrast to Barwise’s $\mathsf{KP}$ in [Bar], which includes the full scheme of foundation, our version of $\mathsf{KP}$ is finitely axiomatizable. Thus $\mathsf{KPR}$ is finitely axiomatizable.
  5. (e)
[…]
  6. (f)
    For $\mathcal{L}\supseteq\mathcal{L}_{\mathrm{set}}$, the dependent choice scheme, denoted $\Pi_{\infty}^{1}$-$\mathsf{DC}(\mathcal{L})$, consists of the universal closure of $\mathcal{L}$-formulae of the following form:
    $\forall x\exists y\ \varphi(x,y,\vec{z})\longrightarrow\forall x\exists f\ \left[f:\omega\rightarrow\mathrm{V},\ f(0)=x,\mathrm{and}\ \forall n\in\omega\ \varphi(f(n),f(n+1),\vec{z})\right].$
    In the presence of $\mathsf{ZF}(\mathrm{\mathcal{L}})\mathrm{,}$ thanks to the Reflection Theorem (Theorem 2.10), $\Pi_{\infty}^{1}$-$\mathsf{DC}\mathrm{(\mathcal{L})}$ is equivalent to the single sentence $\mathsf{DC}$ below:
    $\forall r\ \forall a\ \left[\begin{array}[]{c}\left(\forall x\in a\ \exists y\in a\ \left\langle x,y\right\rangle\in r\right)\longrightarrow\vskip 6.0pt plus 2.0pt minus 2.0pt\\ \left(\forall x\in a\ \exists f\ \left(f:\omega\rightarrow\mathrm{V},\ f(0)=x,\mathrm{and}\ \forall n\in\omega\ \left\langle f(n),f(n+1)\right\rangle\in r\right)\right)\end{array}\right].$
    Note that $\mathsf{DC}$ is provable in $\mathsf{ZFC}$.
@@
  1. (a)
    We follow the convention of using $M$, $M^{\ast},$ $M_{0}$, etc. to denote (respectively) the universes of structures $\mathcal{M}$, $\mathcal{M}^{\ast},$ $\mathcal{M}_{0},$ etc. We denote the membership relation of $\mathcal{M}$ by $\in^{\mathcal{M}}$; thus an $\mathcal{L}_{\mathrm{set}}$-structure $\mathcal{M}$ is of the form $(M,\in^{\mathcal{M}})$.
  2. (b)
    $\mathrm{Ord}^{\mathcal{M}}$ is the class of ordinals in the sense of $\mathcal{M}$, i.e.,
@@
    where $\mathrm{Ord}(x)$ expresses “$x$ is transitive and is well-ordered by $\in$”. More generally, for a formula $\varphi(\vec{x})$ with parameters from $M$, where $\vec{m}=\left(x_{1},\cdot\cdot\cdot,x_{k}\right)$, we write $\varphi^{\mathcal{M}}$ for:
    $\left\{\vec{m}\in M^{k}:\mathcal{M\models\varphi}\left(m_{1},\cdot\cdot\cdot,m_{k}\right)\right\}.$
    A subset $D$ of $M^{k}$ is $\mathcal{M}$-definable if it is of the form $\varphi^{\mathcal{M}}$ for some choice of $\varphi.$ We write $\mathbb{\omega}^{\mathcal{M}}$ for the set of finite ordinals (i.e., natural numbers) of $\mathcal{M}$, and $\mathbb{\omega}$ for the set of finite ordinals in the real world, whose members we refer to as metatheoretic natural numbers. $\mathcal{M}$ is said to be $\omega$-standard if $\left(\mathbb{\omega},\mathbb{\in}\right)^{\mathcal{M}}\cong\left(\mathbb{\omega},\mathbb{\in}\right).$ For $\alpha\in\mathrm{Ord}^{\mathcal{M}}$ we often use $\mathcal{M}_{\alpha}$ to denote the substructure of $\mathcal{M}$ whose universe is
    $M_{\alpha}:=\left\{m\in M:\mathcal{M}\models m\in\mathrm{V}_{\alpha}\right\},$
    where $\mathrm{V}_{\alpha}$ is defined as usual as $\{x:\mathrm{rank}(x)<\alpha\}$, where $\mathrm{rank}(x)=\sup\{\mathrm{rank}(y)+1:y\in x\}.$